# Intervention Efficiency and Perturbation Validation Framework: Capacity-Aware and Robust Clinical Model Selection under the Rashomon Effect

### v1

@@ Appendix D Detailed Setups of Real Data Experiments
To emulate small-cohort deployments and create multiple independent model selection tasks, we partition each dataset into several disjoint subsets, each containing 100 samples, and perform model selection within each subset. For the Breast Cancer dataset ($n=569$), five non-overlapping 100-sample subsets are formed from 500 samples, while the remaining 69 samples are reserved as part of the external test data. For a given subset, the external test set consists of all remaining samples not included in that subset, i.e., n100=469n-100=469 instances for the Breast Cancer dataset. Within each 100-sample subset, we further divide the data into five different train–validation splits, each time using a distinct portion for validation while keeping the rest for training. Repeating this procedure five times reduces the influence of random partitioning within the subset and enables a more reliable assessment of whether one selection method consistently identifies better models across multiple validation scenarios and possibilities. This design focuses on comparing the stability and consistency of model selection performance, rather than evaluating the final predictive accuracy of any individual model.
@@
For each validation fold of each piece of data and each model selection method, we pick one model from the candidate pool using the chosen metric. The selected model is then evaluated on the external test set associated with this piece using the same metric as in validation. This aligns validation with testing and approximates ground truth generalization. We repeat the procedure across pieces and folds, as explained in the last paragraph, and aggregate the testing performance w.r.t the model evaluation metrics, data pieces, and the whole datasets.
@@ Appendix E Sensitivity Analysis
Figure 12: Performance Across Different Noise Added Standard Deviation for Original Size of 50, and $\mu$ ranges from 0.1 to 0.9
@@
Figure 13: Performance Across Different Noise Added Standard Deviation for Original Size of 50, and $\mu$ ranges from 1.1 to 1.9
Figure 14: Performance Across Different Noise Added Standard Deviation for Original Size of 50, and $\mu$ ranges from 2.1 to 2.9
@@
Figure 15: Performance Across Different Noise Added Standard Deviation for Original Size of 100, and $\mu$ ranges from 0.1 to 0.9
@@
Figure 16: Performance Across Different Noise Added Standard Deviation for Original Size of 100, and $\mu$ ranges from 1.1 to 1.9
@@
Figure 17: Performance Across Different Noise Added Standard Deviation for Original Size of 100, and $\mu$ ranges from 2.1 to 2.9
@@
In this subsection, like in the synthetic sensitivity analysis subsection, we isolate the effect of the perturbation noise standard deviation $\sigma$ by looking at the results of two datasets separately and examining PVF’s performance under all possible choices of $\sigma$.
@@
Figure 18: Best Performance for Cerivical Cancer Dataset under Different Evaluation Metrics
@@
Figure 19: Best Performance for Breast Cancer Dataset under Different Evaluation Metrics
@@
The two datasets demand different perturbation scales: small $\sigma$ is most effective for cervical cancer (with a clear optimum near $10^{-2}$ at $\gamma=0.1$), whereas breast cancer benefits from large $\sigma$’s (peaking near 0.30.30.40.4 for smaller capacities and F1 Score) or moderate $\sigma$’s (peaking near $10^{-2}$ for larger capacities), depending on the metric of interest. These patterns underscore that $\sigma$ must be tuned per dataset and evaluation setting rather than being kept as a fixed prior.

### v6

@@ Appendix D Detailed Setups of Real Data Experiments
To emulate small-cohort deployments and create multiple independent model selection tasks, we partition each dataset into several disjoint subsets, each containing 100 samples, and perform model selection within each subset. For the Breast Cancer dataset ($n=569$), five non-overlapping 100-sample subsets are formed from 500 samples, while the remaining 69 samples are reserved as part of the external test data. For a given subset, the external test set consists of all remaining samples not included in that subset, i.e., n100=469n-100=469 instances for the Breast Cancer dataset. Within each 100-sample subset, we further divide the data into five different train–validation splits, each time using a distinct split for validation while keeping the rest for training. Repeating this procedure five times reduces the influence of random partitioning within the subset and enables a more reliable assessment of whether one selection method consistently identifies better models across multiple validation scenarios and possibilities. This design focuses on comparing the stability and consistency of model selection performance, rather than evaluating the final predictive accuracy of any individual model.
@@
For each validation split of each data subset and each model selection method, we pick one model from the candidate pool using the chosen metric based on validation performance. The selected model is then evaluated on the external test set associated with this piece using the same metric as in validation. This aligns validation with testing and approximates ground truth generalization. We repeat the procedure across subsets and splits, as explained in the last paragraph, and aggregate the testing performance w.r.t the model evaluation metrics, subsets and splits.
@@ Appendix E Sensitivity Analysis
Figure 12: The performance difference between PVF and the traditional method on synthetic datasets ($n=50,\mu\in[0.1,0.9]$) across perturbation levels $\sigma$ and evaluation metrics. PVF maintains a stable positive advantage at relatively small $\sigma$ ($\sigma\leq 10^{-2}$), while larger perturbations slightly reduce its margin and can overturn it for F1 Score/accuracy.
Figure 13: The performance difference between PVF and the traditional method on synthetic datasets ($n=50,\mu\in[1.1,1.9]$) across perturbation levels $\sigma$ and evaluation metrics. The advantage of PVF dips as $\sigma$ increases from small to moderate values, but recovers at the largest perturbation level ($\sigma=0.1$).
Figure 14: The performance difference between PVF and the traditional method on synthetic datasets ($n=50,\mu\in[2.1,2.9]$) across perturbation levels $\sigma$ and metrics. Larger perturbations consistently amplify PVF’s advantage, with only minor fluctuations when $\gamma=0.9$.
Figure 15: The performance difference between PVF and the traditional method on synthetic datasets ($n=100,\mu\in[0.1,0.9]$) across perturbation levels $\sigma$ and metrics. PVF maintains a positive margin over small to moderate $\sigma$, while the largest large perturbation reverses the advantage, similarly to the $n=50$ setting.
Figure 16: The performance difference between PVF and the traditional method on synthetic datasets ($n=100,\mu\in[1.1,1.9]$) across perturbation levels $\sigma$ and metrics. PVF shows stable gains at small $\sigma$, dips at mid-range $\sigma$, and recovers at the largest perturbation level, with the pattern varying by evaluation metrics.
Figure 17: The performance difference between PVF and the traditional method on synthetic datasets ($n=100,\mu\in[2.1,2.9]$) across perturbation levels $\sigma$ and metrics. Increasing $\sigma$ steadily amplifies PVF’s advantage, producing especially strong improvements at the largest perturbation level.
@@
In this subsection, like in the synthetic sensitivity analysis subsection, we isolate the effect of the perturbation noise standard deviation $\sigma$ by looking at the results of two datasets separately and examining PVF’s performance under all possible choices of $\sigma$. The performance difference is defined as the proportion of repetitive experiments in which PVF selects a superior model based on the external test set minus the proportion in which the traditional method does, across all subsets and splits.
@@
Figure 18: The performance difference between PVF and the traditional method on the cervical cancer dataset across perturbation levels $\sigma$ and evaluation metrics. PVF achieves the most reliable improvements at small $\sigma$ values (around $10^{-2}$), while larger perturbations diminish or reverse its advantage, especially when $\gamma\geq 0.3$.
@@
Figure 19: The performance difference between PVF and the traditional method on the breast cancer dataset across perturbation levels $\sigma$ and evaluation metrics. PVF benefits most from moderate perturbations ($\sigma\in[10^{-3},10^{-1}]$), particularly for $\gamma\in\{0.1,0.3\}$ and F1 Score, whereas very small or very large $\sigma$ values can lead to ties or reduced gains depending on the metric.
@@
The two datasets demand different perturbation scales: small $\sigma$ is most effective for cervical cancer (with a clear optimum near $10^{-2}$ at $\gamma=0.1$), whereas breast cancer benefits from large $\sigma$’s (peaking near 0.30.30.40.4 for smaller capacities and F1 Score) or moderate $\sigma$’s (peaking near $10^{-2}$ for larger capacities), depending on the metric of interest. These patterns underscore that $\sigma$ should be tuned per dataset and evaluation setting rather than being kept as a fixed prior. With that being said, empirically, we can see that a moderate perturbation (e.g., $\gamma=10^{-2}$) generally applies well to both datasets, which could be a starting point for other similar datasets.